Ed wants to have $1,000,000 at the start of his retirement in 30 years. He currently has $250,000 in his 401K account. If he does not make any more contributions to his account, what annual rate of return must he earn to accomplish his goals?
Answer #1

Ans 4.73

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
1000000 / 250000= (1 + r%)^30
4 = (1 + r%)^30
r = 4.73%